OpenStudy (anonymous):

MUltiply (sqrt6 - 15sqrt3)(2sqrt6 + 6sqrt3) Simplify answer. Type exact answers using radicals as needed.

OpenStudy (anonymous):

= 2(6) + 6sqrt(18) -30sqrt(18) -15(6)(3)

OpenStudy (anonymous):

-258 - 24sqrt(18)

OpenStudy (anonymous):

\[= -258 -24 \sqrt{9} \sqrt{2} = -258 - 72 \sqrt{2}\]

OpenStudy (anonymous):

There are two mistakes in that answer directly above^

OpenStudy (anonymous):

elecengineer, would you please rewrite my original problem as you understood it? Reason: another student told me I was I was writing the problem incorrectly, and I want to make sure I didn't write that one incorrectly.

OpenStudy (anonymous):

The same as the person above. Just the person above made two mistakes.